OpenStudy (anonymous):

3/5 ÷ 2/7 is: 10/21 6/35 1 2 1/10

OpenStudy (anonymous):

@agent0smith @Loser66 Please Help !

OpenStudy (anonymous):

5/6 x 13/20 HELP!!

OpenStudy (agent0smith):

\[\Large \frac{ a }{ b } \div \frac{ b }{ c } = \frac{ a }{ b } \times \frac{ c }{ b } \]

OpenStudy (anonymous):

I got it, it was 2 1/10.. but can you help with this other question?

OpenStudy (anonymous):

@agent0smith :/

OpenStudy (agent0smith):

If you got the first, you should know how to do the second, it's just multiplying fractions.

OpenStudy (anonymous):

yeah I do, thanks for nothing.